# Warm-Starting Collision-Free Model Predictive Control With Object-Centric Diffusion

### v1

@@ V CONCLUSION
While our current conditioning relies on Slot Attention trained on relatively flat, 2D scenes, the framework is designed to naturally extend to fully 3D settings. Future work will explore richer object-centric representations to further improve generalization. Another exciting direction is an online variant that integrates diffusion directly within the MPC loop and adaptively selects between the previous plan and the newly inferred one, enabling even faster and more responsive control and providing a way to enforce safety guarantees on diffusion policies, which can be integrated with large vision–language action models such as $\pi_{0.5}$[intelligence_05_2025]. We believe these extensions will significantly broaden the applicability of our approach to more complex, dynamic, and cluttered environments.

### v2

@@ V CONCLUSION
While our current conditioning relies on Slot Attention trained on relatively flat, 2D scenes, the framework is designed to naturally extend to fully 3D settings. Future work will explore richer object-centric representations to further improve generalization. Another exciting direction is an online variant that integrates diffusion directly within the MPC loop and adaptively selects between the previous plan and the newly inferred one, enabling even faster and more responsive control and providing a way to enforce safety guarantees on diffusion policies, which can be integrated with large vision–language action models such as $\pi_{0.5}$[49]. We believe these extensions will significantly broaden the applicability of our approach to more complex, dynamic, and cluttered environments.